Company XYZ has a total debt to total assets ratio of 40%, what is company’s equity multiplier?
Answer #1

Total Debt to Total Assets = 40%

Equity multiplier = Total Asset/Equity = 1/(1-Total Debt/Total Assets ) = 1/(1-40%) = 100/60 = 5/3 = 1.67

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well